# Asymmetric Duos: Sidekicks Improve Uncertainty

### v1

@@ 5 Experiment Results
Here we present the performance of Asymmetric Duos, as measured by accuracy, uncertainty quantification, and selective classification performance, as a function of computation (i.e. Duo balance). For each benchmark, we form Asymmetric Duos using all possible combinations of $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ architectures outlined in Section 4. For visual clarity, this section only contains results for a single $f_{\mathrm{large}}$ on each dataset. (We choose different $f_{\mathrm{large}}$ architectures for each dataset for diversity of results within this section.) We note that the results in this section are representative of all $f_{\mathrm{large}}$ architectures (see Appendix E for the full set of results).
@@ 5.1 Duos as Classifiers – Accuracy & Macro F1 Score
Figure 3: Class Prediction (Accuracy and F1 $\uparrow$) as a function of FLOPs balance. ($\text{Balance}=0$ corresponds to cost of a single $f_{\mathrm{large}}$ model; $1$ corresponds to the cost of a $m=2$ ensemble.) Asymmetric Duos almost always increase accuracy over a single $f_{\mathrm{large}}$, even for Duos that combine $f_{\mathrm{large}}$ with a $f_{\mathrm{small}}$ that is $1/10^{\mathrm{th}}$ the size. The learned temperature weighting is crucial to achieve this performance; if the predictions of $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ are averaged with equally weight (Duo: Unweighted) then imbalanced Duos may have significantly worse performance than single models.
@@
In Appendix E we also report results for probabilistic measures of accuracy; namely Brier Score and negative log likelihood. The results for these metrics follow similar trends.
@@ 5.2 Predictive Uncertainty Separability - Correctness Prediction AUROC
Predictive uncertainty estimates are often judged by their calibration, as measured by Expected Calibration Error (ECE) [e.g. 20, 41]. Because we apply temperature scaling throughout our experiments, the $f_{\mathrm{large}}$ models are largely calibrated to begin with, and so Asymmetric Duos provide little improvement. (See Appendix E for results.) We argue that this is not a failure of Duos but rather a success of temperature scaling.
We instead choose to measure the quality of Duo uncertainty estimates based on their correlation with the correctness of class predictions. Intuitively, a good uncertainty estimate should be able to separate corerct versus incorrect class predictions. To that end, we report correctness prediction Area Under the Receiver Operating Characteristic curve (AUROC) for Duos and for single $f_{\mathrm{large}}$ as a function of FLOPs balance. Randomly assigned uncertainty measure would yield an AUROC score of $0.5$, while a perfectly separable uncertainty measure would yield $1.0$. We note that the AUROC metric couples accuracy and uncertainty quantification, in that any improvement in AUROC could be attributed either to changes in accuracy or improvements in uncertainty separability.
In Figure 4 we find that Asymmetric Duos always yield higher AUROC than $f_{\mathrm{large}}$. Surprisingly, we observe performance improvements even when the FLOPs balance is near 00. This finding suggests that AUROC can be increased without much additional computation. Performance does not correlate with additional computation (i.e. higher FLOPs balance), but all Duos provide benefit.
Our ablation demonstrates that these AUROC improvements cannot solely be attributed to the increased accuracy afforded by Duos. Recall that the UQ Only Duo variant uses the Duo’s uncertainty estimate while retaining the class prediction of $f_{\mathrm{large}}$. Any AUROC benefits must therefore be attributed to improvements in the uncertainty, as accuracy remains unchanged. We observe that UQ Only Duos obtain significantly better AUROC score compared to their base models, confirming that Duos provide meaningful improvements in uncertainty separability in addition to accuracy.
Figure 4: Correctness prediction as measured by AUROC ($\uparrow$), which captures the separability of correct and incorrect predictions by uncertainty. In almost all cases, Asymmetric Duos achieve higher AUROC than corresponding $f_{\mathrm{large}}$ models in isolation, even when using a neglegable amount of additional computation. We confirm that this increase cannot be attributed to accuracy improvements alone: our ablation (Duo: UQ only) that uses the Duo’s uncertainty measure to separate the $f_{\mathrm{large}}$ class prediction produces a comparable AUROC increase.
@@ 5.3 Duos as selective classifiers – AURC
To study the effect of Duo uncertainty estimates on downstream decision-making tasks, we evaluate Asymmetric Duos in the context of selective classification. Selective classification allows a model to abstain from prediction when the uncertainty is high, deferring uncertain cases to experts in effective human-in-the-loop systems [15]. The proportion of the test set with uncertainty below a given threshold is referred to as coverage, and performance on this subset is measured using the domain-specific loss, which in our case is the 0-1 loss analogous to classification error rate. As the uncertainty threshold decreases, the model becomes more cautious, abstaining from more inputs and reducing coverage while aiming to improve accuracy on this smaller covered subset. We emphasize that high-quality uncertainty estimates are necessary for selective classifiers to work in practice.
We evaluate selective classification performance using the Area Under the Risk Coverage (AURC) metric [16]. AURC accumulates error rates at all coverage levels to measure a model’s overall potential as a selective classifier. As with correctness prediction, the AURC metric conflates accuracy and uncertainty quantification, as different predictors achieve different accuracy at full coverage (entire test set accuracy). To that end, we evaluate our UQ Only Duo ablation in addition to the proposed approach to bypass this potential confound.
@@
In Figure 5, we observe that Duos provide subtantial boost to AURC at all computation levels, even when the FLOPs balance is fractional. Analoguos to our correctness prediction results, UQ Only Duos also achieve improved AURC scores at low FLOPs balance, demonstrating that these improvements cannot be attributed to accuracy alone. This is a direct result of our findings on correctness prediction with AUROC. Together, these results also confirm that smaller models can effectively re-rank the uncertainties of larger models. In contrast, many prior methods for selective classification sacrifice non-selective predictive power [13], resulting in unfavorable trade-offs in practical metrics such as AURC.
@@ 5.4 Duos as selective classifiers – Selective Accuracy Constraint
To further evaluate Duos in the selective classification context, we report the Selective Accuracy Constraint (SAC) metric for all Duos and $f_{\mathrm{large}}$ models. SAC is a practical evaluation metric for risk-sensitive applications with a guarantee. It measures the maximum coverage a predictor can achieve while maintaining a fixed target accuracy, making it well-suited for settings where reliability is critical. Prior work in selective classification has shown that leveraging softmax response uncertainty allows deep networks to achieve high accuracy on a subset of confidently predicted data by abstaining on uncertain cases [15, 13]. Here we show duos boost coverage at the ambitious accuracy levels of 98% across all of our benchmarked datasets.
@@ 5.5 Soup with a (Side)kick: Combining Model Soups and Duos
We show the Asymmetric Duo framework is agnostic to model type, compatible with other methods, and compounds uncertainty quantification improvements by making a duo with a model soup [58]. For this experiment, we fine-tuned thirty-two ConvNeXt Base models on Caltech 256 starting from the same torchvision pre-trained model using different training hyperparameters. We then mixed a greedy model soup for use as $f_{\mathrm{large}}$ in our Asymmetric Duo. In Figure 7, we observe that Duos are compatible with soups and other methods of making better base and sidekick models.
@@ 6 Discussion
Gontijo-Lopes et al. [18] used a similar construction of heterogeneous $m=2$ deep ensembles to study the effect of predictive diversity induced by models pre-trained on different datasets. Our work differs mainly by aggregating deep network pairs of vastly different sizes as opposed to different architectures with similar accuracy, and diving deeper into a smaller member model’s effects on a Duo’s UQ capability.
Rahaman and Thiery [46] studied the combination of temperature scaling [20] and deep ensembles [28] and found that deep ensembles constructed by pool-then-calibrate or pool-with-calibrate strategies outperform naive deep ensembles and those following a calibrate-then-pool strategy on proper scoring rules. Our proposed Asymmetric Duos follow a similar construction order to a pool-with-calibration strategy, but fundamentally differ in motivation. As shown with our Unweighted Duos ablation, the asymmetry between $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ makes a pool-with-calibrate strategy necessary in preventing degradation in accuracy. Whereas Rahaman and Thiery [46] focuses on such constructions’ effects on calibration, we put a unique emphasis on uncertainty quantification metrics like correctness prediction and selective classification.
@@
This work serves as an initial step toward understanding the aggregation strategies of deep networks of different sizes and capacities. While we focused on a small set of interpretable aggregation strategies for both point prediction and uncertainty quantification, there remains a large space of unexplored methods. Future work may benefit from exploring other learning-based, adaptive weighting schemes for member aggregation, and may even call them Dynamic Duos.
@@ A.1 Model Architectures
We select popular, recent, and performative pre-trained models from torchvision [36] and timm [57] to construct our Duos, to make sure Duos are compatible with various recent architectures. From torchvision, we experiment with EfficientNet [50], EfficientNet V2 [51], MNASNet [52], ConvNeXt [34], ResNet [23], ResNeXt [59], RegNet [45], ShuffleNet-V2, Swin [60], and Swin (V1 & V2) [33, 32]. From timm, we get an Efficientnet-V2-L pre-trained on ImageNet 21k [7], and a ViT-B-16 [9] pre-trained on LAION 2B [49].
@@ A.2 Training Procedure
For all our fine-tuned experiments we follow the LP–FT recipe from Kumar et al. [27]: we first train only the final classification head (also known as Linear Probing, or LP) for a few epochs (8 on Caltech 256, 5 on iWildCam) using cross‑entropy loss and AdamW, then unfreeze the entire network and Fine‑Tune for more epochs (20 on Caltech 256, 10 on iWildCam) under the same loss and optimizer.
Our fine‑tuning (FT) procedure employs a sequential schedule (linear warm‑up then cosine annealing), and we randomly search learning rate in $[1\times 10^{-6},3\times 10^{-4}]$ and weight decay in $[1\times 10^{-8},1\times 10^{-5}]$, picking the best trial by validation score. We use an adaptive batch‑sizing strategy that finds the optimal (maximum) batch-size for each architecture allowed by training resources, capped at 128. Random Augmentation [5] is used to augment training samples during the FT phase. The best LP checkpoint initializes FT, and the top FT model by validation performance is carried forward into our Duo evaluations.
@@ A.3 Datasets
ImageNet [7] contains 1,000 classes and more than a million training samples. All the pretrained models used in our experiment have been trained on ImageNet as a last step. Since ImageNet never released the official test set, researchers split the official validation set into val and test splits. Since we don’t need to fine-tune models on ImageNet for our experiment, our only usage of a validation set is to temperature scale single models and to tune temperature weightings for Asymmetric Duos. We use only 5% of the official test split as our validation set to show how data-efficient the temperature-weighting step is.
ImageNet V2 [47] is a careful reproduction of the original ImageNet, but nevertheless showed model performance degradation, indicating distribution shifts. Since this dataset is used solely for evaluation, the entire dataset is treated as an OOD test set for ImageNet models and Duos.
iWildCam (OOD) [3, 26] is a camera-trap dataset with 203,029 training images containing 182 classes of animals. The IND and OOD are determined by distinctive camera traps, since different locations vary largely in vegetation and backgrounds, resulting in poor generalization to new camera traps. We use the official IND val split for iWildCam, which contains 7315 images captured by the same set of camera traps as the training set.
Caltech 256 [19] contains 30,607 images spanning 256 object classes and a clutter class. We use a random 15% split of the dataset as val and another 15% as test. The same validation set was used for both hyper-tuning and temperature tuning.
A.4 Greedy Soup
For our Greedy Soup experiments, we independently train 32 ConvNeXt-Base models initialized with different hyperparameters over the same search space as our regular tuning procedure on Caltech 256 starting from the same pre-trained weights from torchvision, and greedily average their weights to improve performance on the validation set following the procedure outlined by Wortsman et al. [58]. The greedy soup in Figure 7 averages five of the trials.

### v2

@@ 5 Experiment Results
Here we present the performance of Asymmetric Duos, as measured by accuracy, uncertainty quantification, and selective classification performance, as a function of computation (i.e. Duo FLOPs balance). For each benchmark, we form Asymmetric Duos using all possible combinations of $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ architectures outlined in Section 4. For visual clarity, this section only contains results for a single $f_{\mathrm{large}}$ on each dataset. (We choose different $f_{\mathrm{large}}$ architectures for each dataset for diversity of results within this section.) We note that the results in this section are representative of all $f_{\mathrm{large}}$ architectures (see Appendix G for the full set of results).
@@ 5.1 Duos as Classifiers – Accuracy & Macro F1 Score
Figure 3: Class Prediction (Accuracy and F1 $\uparrow$) as a function of FLOPs balance. ($\text{Balance}=0$ corresponds to cost of a single $f_{\mathrm{large}}$ model; 11 corresponds to the cost of a $m=2$ ensemble.) Asymmetric Duos almost always increase accuracy over a single $f_{\mathrm{large}}$, even for Duos that combine $f_{\mathrm{large}}$ with a $f_{\mathrm{small}}$ that is $1/10^{\mathrm{th}}$ the size. The learned temperature weighting is crucial to achieve this performance; if the predictions of $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ are averaged with equally weight (Duo: Unweighted) then imbalanced Duos may have significantly worse performance than single models.
@@
In Appendix G we also report results for probabilistic measures of accuracy; namely Brier Score and negative log likelihood. The results for these metrics follow similar trends.
@@ 5.2 Predictive Uncertainty Separability - Correctness Prediction AUROC
Predictive uncertainty estimates are often judged by their calibration, as measured by Expected Calibration Error (ECE) [e.g. guo2017calibration, ovadia2019can]. Because we apply temperature scaling throughout our experiments, the $f_{\mathrm{large}}$ models are largely calibrated to begin with, and so Asymmetric Duos provide little improvement. (See Appendix G for results.) We argue that this is not a failure of Duos but rather a success of temperature scaling.
We instead choose to measure the quality of Duo uncertainty estimates based on their correlation with the correctness of class predictions. Intuitively, a good uncertainty estimate should be able to separate correct versus incorrect class predictions. To that end, we report correctness prediction Area Under the Receiver Operating Characteristic curve (AUROC) for Duos and for single $f_{\mathrm{large}}$ as a function of FLOPs balance. Randomly assigned uncertainty measure would yield an AUROC score of $0.5$, while a perfectly separable uncertainty measure would yield $1.0$.
In Figure 4 we find that Asymmetric Duos consistently yield higher AUROC than $f_{\mathrm{large}}$. Surprisingly, we observe performance improvements even when the FLOPs balance is near 0. This finding suggests that AUROC can be increased without much additional computation. Performance does not correlate with additional computation (i.e. higher FLOPs balance), but all Duos provide benefit.
Our ablation demonstrates that these AUROC improvements cannot solely be attributed to the increased accuracy afforded by Duos. Recall that the UQ Only Duo variant uses the Duo’s uncertainty estimate while retaining the class prediction of $f_{\mathrm{large}}$. Any AUROC benefits must therefore be attributed to improvements in the uncertainty, as accuracy remains unchanged. We observe that UQ Only Duos obtain significantly better AUROC score compared to their base models, showing that the sidekick model is effective at re-ranking uncertainties for the class prediction of $f_{\mathrm{large}}$, confirming that Duos provide meaningful improvements in uncertainty separability in addition to accuracy.
Figure 4: Correctness prediction as measured by AUROC ($\uparrow$), which captures the separability of correct and incorrect predictions by uncertainty. In almost all cases, Asymmetric Duos achieve higher AUROC than corresponding $f_{\mathrm{large}}$ models in isolation, even when using a negligible amount of additional computation. We confirm that this increase cannot be attributed to accuracy improvements alone: our ablation (Duo: UQ only) that uses the Duo’s uncertainty measure to separate the $f_{\mathrm{large}}$ class prediction produces a comparable AUROC increase.
@@ 5.3 Duos as selective classifiers – AURC
To study the effect of Duo uncertainty estimates on downstream decision-making tasks, we evaluate Asymmetric Duos in the context of selective classification. Selective classification allows a model to abstain from prediction when the uncertainty is high, deferring uncertain cases to experts in effective human-in-the-loop systems [geifman2017selective]. The proportion of the test set with uncertainty below a given threshold is referred to as coverage, and performance on this subset is measured using the domain-specific loss, which in our case is the 0-1 loss analogous to classification error rate. As the uncertainty threshold decreases, the model becomes more cautious, abstaining from more inputs and reducing coverage while aiming to improve accuracy on this smaller covered subset. We emphasize that high-quality uncertainty estimates are necessary for selective classifiers to work in practice.
We evaluate selective classification performance using the Area Under the Risk Coverage (AURC) metric [geifman2018biasreduced]. AURC accumulates error rates at all coverage levels to measure a model’s overall potential as a selective classifier. As with correctness prediction, the AURC metric conflates accuracy and uncertainty quantification, as different predictors achieve different accuracy at full coverage (entire test set accuracy). To that end, we evaluate our UQ Only Duo ablation in addition to the proposed approach to bypass this potential confound.
@@
In Figure 5, we observe that Duos provide substantial boost to AURC at all computation levels, even when the FLOPs balance is fractional. Analogous to our correctness prediction results, UQ Only Duos also achieve improved AURC scores at low FLOPs balance, demonstrating that these improvements cannot be attributed to accuracy alone. This is a direct result of our findings on correctness prediction with AUROC. Together, these results also confirm that smaller models can effectively re-rank the uncertainties of larger models. In contrast, many prior methods for selective classification sacrifice non-selective predictive power [galil2023what], resulting in unfavorable trade-offs in practical metrics such as AURC.
@@ 5.4 Duos as selective classifiers – Selective Accuracy Constraint
To further evaluate Duos in the selective classification context, we report the Selective Accuracy Constraint (SAC) metric for all Duos and $f_{\mathrm{large}}$ models. SAC is a practical evaluation metric for risk-sensitive applications with a guarantee. It measures the maximum coverage a predictor can achieve while maintaining a fixed target accuracy, making it well-suited for settings where reliability is critical. Prior work in selective classification has shown that leveraging softmax response uncertainty allows deep networks to achieve high accuracy on a subset of confidently predicted data by abstaining on uncertain cases [geifman2017selective, galil2023what]. Here we show duos boost coverage at the ambitious accuracy levels of 98% across all of our benchmarked datasets.
@@ 5.5 Soup with a (Side)kick: Combining Model Soups and Duos
We show the Asymmetric Duo framework is agnostic to model type, compatible with other methods, and compounds uncertainty quantification improvements by making a duo with a model soup [wortsman2022model]. For this experiment, we fine-tuned sixteen ConvNeXt Base models on Caltech 256 starting from the same torchvision pre-trained model using different training hyperparameters. We then mixed a greedy model soup for use as $f_{\mathrm{large}}$ in our Asymmetric Duo. In Figure 7, we observe that Duos are compatible with soups and other methods of making better base and sidekick models.
@@
5.6 Comparisons between Asymmetric Duos and Deep Ensembles
We show that Asymmetric Duos (AD) can match the performance of classic deep ensembles (DE) with a significantly lower compute budget. We construct ensembles of ResNet50s (RN50) trained on the ImageNet dataset following the “pool-then-calibrate” strategy from [rahaman2021uncertainty]. We compare against ADs that pair a RN50 ($f_{\mathrm{large}}$) with MNASNet0.75 (MN.75), ShuffleNet V2 2.0× (SN2.0x), and EfficientNet-B2 (ENB2) models, yielding FLOPs Balances (FB) of 0.05, 0.14, and 0.27, respectively.
Table 1: Base-model-controlled comparison between Asymmetric Duos and deep ensembles.
[…]
Table 1 shows that Asymmetric Duos quickly approach performance of $m=5$ Deep Ensembles (e.g., $m=5$ RN50s) across most metrics, despite using $4\times$ fewer FLOPs. We attribute this to architectural diversity from heterogeneity [gontijo-lopes2022no] and Duos’ simple but effective temperature-based aggregation. If we allow Asymmetric Duos to match the computational budget of $m=5$ Deep Ensembles, then Duos become even more advantageous. See Appendix E for a FLOPs-controlled comparison.
@@ 6 Discussion
gontijo-lopes2022no used a similar construction of heterogeneous $m=2$ deep ensembles to study the effect of predictive diversity induced by models pre-trained on different datasets. Our work differs mainly by aggregating deep network pairs of vastly different sizes as opposed to different architectures with similar accuracy, diving deeper into a smaller member model’s effects on a Duo’s uncertainty, and of course our fine-tuning the pairs on the same data.
rahaman2021uncertainty studied the combination of temperature scaling [guo2017calibration] and deep ensembles [lakshminarayanan2017simple] and found that deep ensembles constructed by pool-then-calibrate or pool-with-calibrate strategies outperform naive deep ensembles and those following a calibrate-then-pool strategy on proper scoring rules. Our proposed Asymmetric Duos follow a similar construction order to a pool-with-calibration strategy, but fundamentally differ in motivation. As shown with our Unweighted Duos ablation, the asymmetry between $f_{\mathrm{large}}$ and $f_{\mathrm{small}}$ makes a pool-with-calibrate strategy necessary in preventing degradation in accuracy. Whereas rahaman2021uncertainty focuses on such constructions’ effects on calibration, we put a unique emphasis on uncertainty quantification metrics like correctness prediction and selective classification.
@@
This work serves as an initial step toward understanding the aggregation strategies of deep networks of different sizes and capacities. While we focused on a small set of interpretable aggregation strategies for both point prediction and uncertainty quantification, there remains a large space of unexplored methods. Notably, our finding that simple validation-based weighting can yield effective ensembles from models with vastly different accuracies is itself a surprising and underexplored result. Future work may benefit from exploring other learning-based, adaptive weighting schemes for member aggregation, and may even call them Dynamic Duos.
@@ A.1 Model Architectures
We select popular, recent, and performative pre-trained models from torchvision [torchvision2016] and timm [rw2019timm] to construct our Duos, to make sure Duos are compatible with various recent architectures. From torchvision, we experiment with EfficientNet [tan2019efficientnet], EfficientNet V2 [tan2021efficientnetv2], MNASNet [tan2019mnasnet], ConvNeXt [liu2022convnet], ResNet [he2016deep], ResNeXt [xie2017aggregated], RegNet [radosavovic2020designing], ShuffleNet-V2, Swin [zhang2018shufflenet], and Swin (V1 & V2) [liu2022swin, liu2021swin]. From timm, we get an Efficientnet-V2-L pre-trained on ImageNet 21k [deng2009imagenet], and a ViT-B-16 [dosovitskiy2021an] pre-trained on LAION 2B [schuhmann2022laionb].
@@ A.2 Training Procedure
For all our fine-tuned experiments we follow the LP–FT recipe from kumar2022fine: we first train only the final classification head (also known as Linear Probing, or LP) for a few epochs (8 for both Caltech 256 and iWildCam) using cross‑entropy loss and AdamW, then unfreeze the entire network and Fine‑Tune (FT) for more epochs (16 on Caltech 256, 12 on iWildCam) under the same loss and optimizer.
Our FT procedure employs a sequential schedule (linear warm‑up followed by cosine annealing), and we perform hyper-parameter search over learning rate in $1\times 10^{-6},3\times 10^{-4}]$ and weight decay in $1\times 10^{-8},1\times 10^{-5}]$, picking the best trial by validation score. We use a batch size of 128 for Caltech 256 and 16 for iWildCam. All models are trained on NVIDIA L40S GPUs. Random Augmentation [cubuk2020randaugment] is used to augment training samples during the FT phase. The best LP checkpoint initializes FT, and the top FT model by validation performance is carried forward into our Duo evaluations.
@@ A.3 Datasets
ImageNet [deng2009imagenet] contains 1,000 classes and over a million training samples. All the pretrained models used in our experiment have been trained on ImageNet as a last step. Since ImageNet never released the official test set, researchers split the official validation set into val and test splits. Since we don’t need to fine-tune models on ImageNet for our experiment, our only usage of a validation set is to temperature scale single models and to tune temperature weightings for Asymmetric Duos. We use only 5% of the official test split as our validation set to show how data-efficient the temperature-weighting step is.
ImageNet V2 [recht2019imagenet] is a careful reproduction of the original ImageNet, but nevertheless showed model performance degradation, indicating distribution shifts. Since this dataset is used solely for evaluation, the entire dataset is treated as an OOD test set for ImageNet models and Duos.
iWildCam (OOD) [beery2020iwildcam2020competitiondataset, koh2021wilds] is a camera-trap dataset with 203,029 training images containing 182 classes of animals. The IND and OOD are determined by distinctive camera traps, since different locations vary largely in vegetation and backgrounds, resulting in poor generalization to new camera traps. We use the official IND val split for iWildCam, which contains 7315 images captured by the same set of camera traps as the training set.
Caltech 256 [griffin2007caltech] contains 30,607 images spanning 256 object classes and a clutter class. We use a random 15% split of the dataset as val and another 15% as test. The same validation set was used for both hyper-tuning and temperature tuning.
A.4 Greedy Soup Setup
For our Greedy Soup experiments, we independently fine-tune 16 ConvNeXt-Base models initialized from the same backbone with different hyperparameters over the same search space as our regular tuning procedure on Caltech 256 starting from the same pre-trained weights from torchvision, and greedily average their weights to improve performance on the validation set following the procedure outlined by wortsman2022model. The greedy soup in Figure 7 averages three of the sixteen trials.
A.5 Deep Ensemble Setup
All $m=x$ DEs in Section 5.6 and Appendix E are constructed by randomly selecting 32 subsets of size $x$ from 10 independently trained ResNet-50 checkpoints from timm and torchvision, following the “pool-then-calibrate” strategy of rahaman2021uncertainty. Because our temperature-based aggregation strategy for Duos naturally incorporates calibration, this setup enables a fair comparison on NLL, Brier, and ECE metrics.
All ADs used for comparison in Section 5.6 are formed by pairing one ResNet-50 checkpoint as $f_{\mathrm{large}}$ with a sidekick $f_{\mathrm{small}}$ drawn from torchvision.